OpenStudy (cucacula):

calculate the coordinates of the critical path points of the curve y = x^3 - 6x^2 and sketch the graph . anyone can help me?

OpenStudy (anonymous):

critical points -> points where the "derivative" becomes 0 so, first differentiate the function.

OpenStudy (cucacula):

can you show me the 1st step? @electrokid

OpenStudy (anonymous):

\[y=x^3-6x^2\\ {dy\over dx}=y'={d\over dx}(x^3-6x^2)=?\]

OpenStudy (cucacula):

isit d/dx ( 3x^2 - 12x ) @electrokid

OpenStudy (anonymous):

good.. without the d/dx part :) since you already evaluated it,

OpenStudy (anonymous):

now, set it equal to zero and solve for ":x"

OpenStudy (cucacula):

can give me the hint? @electrokid

OpenStudy (anonymous):

\[3x^2-12x=0\] solve for "x"

OpenStudy (cucacula):

3x^2 = 12x x^2 = 12x/3 x^2 = 4x x = 4 isit true? @electrokid

OpenStudy (anonymous):

no.. you missed one possible solution... \[3x^2-12x=0\\ 3x(x-4)=0\\\boxed{x=0}\qquad{\rm OR}\qquad\boxed{x=4}\]

OpenStudy (cucacula):

so the x = 0 or x = 4 ? @electrokid can u sketch the graph?

OpenStudy (anonymous):

those are the critical points.